OpenStudy (anonymous):

How do I find the x intercept and y intercept by the function f(x) = 400 - 4/5x?

hero (hero):

To find the x intercept, set y = 0, then solve for x: Remember y = f(x) so you have y = 400 - (4/5)x Set y = 0 0 = 400 - (4/5)x Add (4/5)x to both sides: (4/5)x = 400 Multiply both sides by (5/4) x = (5/4)(400) x = (5)(100) x = 500

hero (hero):

To find the y-intercept, set x = 0, then solve for y.